OpenStudy (kennyval):

Fifteen more than four times a number is 6 more than five times the number. Use "n" for the number.

OpenStudy (soggylearnsandhelps):

This seems hard.

OpenStudy (soggylearnsandhelps):

Let a number be:x "15 more than" means: +15 "4 times a number" means: 4x "is" means:= "6 more than" mean:+6 "5 times the number" means: 5x Problem to solve: 4x+15=5x+6 -4x+4x+15=-4x+5x+6 15=x+6 15-6=x+6-6 9=x A number is: x=9 (e)

OpenStudy (soggylearnsandhelps):

Ah you beat me!

zepdrix (zepdrix):

lol you said it was hard :P didn't think you were going to write anything XD

OpenStudy (davejavous):

\[4n+15=5n+6\] \[15−6=5n−4n\] \[9=n\] \[n=9\]